OpenStudy (anonymous):

I'm thinking of a number from 1 to 10. It's prime. Twice the number plus one is prime. Twice the number minus one is not prime. What's the number?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

is 1 a prime or 3*2 + 1 = 7

OpenStudy (anonymous):

5

OpenStudy (anonymous):

umm no it doesn't work

OpenStudy (anonymous):

2 does not work

OpenStudy (anonymous):

is 1 a prime??

OpenStudy (anonymous):

1 is not prime

OpenStudy (anonymous):

no, the first prime is 2

OpenStudy (anonymous):

2,3,5,7

OpenStudy (mimi_x3):

the number is 3?

OpenStudy (anonymous):

nope, that fails "twice the number minus one is not prime"

OpenStudy (anonymous):

it's 5

OpenStudy (anonymous):

i got it iknow it has to do something with numbers humm 69?

OpenStudy (anonymous):

oh yeah 5

OpenStudy (anonymous):

5*2+1=11 which is prime 5*2-1=9 which is not prime

OpenStudy (anonymous):

ghass1978 was right

OpenStudy (anonymous):

ding ding ding

OpenStudy (anonymous):

hmm how could i fall to such a question ughhh

OpenStudy (mimi_x3):

lols, im feel so stupid, got an easy question wrong xD

OpenStudy (mimi_x3):

i feel*

OpenStudy (anonymous):

{{2, 5, 3}, {3, 7, 5}, {5, 11, 9}, {7, 15, 13}} 5 only.